# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 7.1. Datasets
Check-worthy claim detection is a popular task within the NLP community thanks to the series of CheckThat! shared tasks organised at CLEF, which led to the release of related datasets. Indeed, check-worthy claim detection is the only subtask that has been proposed at all seven CheckThat! editions (Atanasova et al., 2018, 2019; Barrón-Cedeño et al., 2020; Nakov et al., 2021a, 2022a; Alam et al., 2023; Barrón-Cedeño et al., 2024). Through those editions, datasets for training check-worthy claim detection models have been constantly extended to cover additional languages, starting from English and Arabic at CheckThat! 2018 (Atanasova et al., 2018) to Arabic, Dutch, English and Spanish at CheckThat! 2024 (Barrón-Cedeño et al., 2024) and even multimodal and multi-genre data in the 2023 edition (Alam et al., 2023). Beside CheckThat!, additional datasets have been created and made available for research, as shown in Table 8. The sources covered by such datasets are typically social media, news and political debates, i.e. three relevant areas where the presence of disinformation may have detrimental effects on a large audience. Particular relevance was given to COVID-related content during the pandemic (Hadj Ameur and Aliane, 2021; Salek Faramarzi et al., 2023), when the continuous flow of misleading information could negatively affect public health. Although English is the most represented language, datasets for check-worthiness detection have been created in several other languages, probably because being relevant and of interest to the general public is something that is time- and geographically-bounded, with claims referring to specific events for each specific country or area.
Dataset | Language | # Claims | Content type | Classes | ||||||
@@
Dataset | Language(s) | # Input claims | # Verified claims | # Pairs | Content type |
@@ 7.2. Methods and models
Existing works try to first of all define what check-worthiness is. In a recent survey, Panchendrarajan and Zubiaga (Panchendrarajan and Zubiaga, 2024) identify two main aspects making a claim check-worthy: its verifiability and its priority. The first item refers to the possibility of determining the veracity of a claim, which can be likely supported by evidence. A claim is verifiable when it contains a factual statement that can be checked, which means that personal opinions or events presented as uncertain are excluded. Second, given that verifying all statements about the world is impossible, it is important to prioritize claims which are considered timely, interesting to the general public and whose verification might have a broader impact (Das et al., 2023). In this latter aspect it differs from a related, but a distinct task of claim detection which solely aims to identify what constitutes a claim in a text (either using a binary classification or by identification of spans) (Mittal et al., 2023; Wuehrl et al., 2023; Gupta et al., 2021b).
@@
Given this need for prioritization, check-worthiness detection has been cast as a ranking problem since the first editions of the CheckThat! Lab within the CLEF Evaluation initiative,242424https://www.clef-initiative.eu/ which made the task well-known within the NLP community, although some related works had already been presented before (Hassan et al., 2015). In particular, given a political debate, the first CheckThat! task was aimed at predicting which claims should be prioritized for fact checking. This is reflected in the evaluation approach proposed for the CheckThat! series, which has become the de facto standard in check-worthiness detection: systems should output the list of input claims ranked by check-worthiness, which is usually evaluated using Mean Average Precision (MAP), reciprocal rank and for . More recently, however, in the shared task check-worthiness has been evaluated using F1 in a binary classification task rather than ranking (Alam et al., 2023; Barrón-Cedeño et al., 2024).
@@
Regarding methods, in the CheckThat! editions up to 2023 state-of-the-art results for check-worthy claim detection were reached by methods that rely on fine-tuned transformer-based methods such as BERT, RoBERTa, DistilBERT, (Williams et al., 2021; Frick et al., 2023; Sawinski et al., 2023) and language-specific variants, (Alam et al., 2023), often combined with data manipulation or model ensembling strategies. For example, the best results on the English portion of the CheckThat 2022 dataset were achieved by a RoBERTa model that leveraged a back-translation-driven data augmentation process (Savchev, 2022). The last edition in 2024, instead, has seen an increased interest in using also generative LLMs for the task such as LLama 2 and 3, GPT3.4 and 5, Mixtral and Mistral (Barrón-Cedeño et al., 2024). For instance, the best performing system on English fine-tuned Llama2 7b on the original training data, using prompts generated by ChatGPT (Li et al., 2024b). Nevertheless, despite the notable progress in data and methods for check-worthy claim detection, there is currently a lack of an extended coverage across languages and topics.
@@ 7.3. Tools and services
ClaimBuster (Hassan et al., 2017) was the first end-to-end system for computer-assisted fact-checking trained on a human-labeled dataset of check-worthy factual claims from the U.S. general election debate transcripts. The first component of the pipeline is a detector of check-worthy factual claims which, given a sentence, first labels it as being ‘non-factual’, ‘factual and unimportant’ or ‘factual and check-worthy’. In case of the latter, a ranking score is assigned based on SVM decision function. Patwari et al. (Patwari et al., 2017) present Tathya, a tool focusing only on check-worthiness detection, which compared to ClaimBuster can yield a significant performance improvement, particularly on recall. It is based on a multi-classifier system using features such as topics, entity history and PoS tuples.
ClaimRank (Jaradat et al., 2018) performs check-worthy claim detection and supports English and Arabic texts. Its strength is that it was trained on actual annotations from nine reputable fact-checking organizations, therefore mimicking their real strategy for claim selection. The ranking is based on a number of lexical, structural and semantic features, used to train a neural network with two hidden layers as proposed by Gencheva et al. (Gencheva et al., 2017). Another system, focusing specifically on tweets in Arabic, is Tahaqqaq (Sheikh Ali et al., 2023), which includes the possibility to identify check-worthy claims, estimate the user credibility in terms of spreading fake news, and find authoritative accounts.
@@
Concerning retrieval of previously fact-checked claims, it is supported by some of the end-to-end verification systems mentioned above, namely ClaimBuster (Hassan et al., 2017), Tahaqqaq (Sheikh Ali et al., 2023) or BRENDA (Botnevik et al., 2020). Besides these, Google Fact-Check Explorer252525https://toolbox.google.com/factcheck/explorer is often used to perform the task since it indexes a large corpus of fact-checks. Other specialized tools include Fact-Check Finder262626https://fact-check-finder.kinit.sk built on models developed in (Pikuliak et al., 2023).
@@ 7.4. Discussion
Multimodality. Although most available datasets are mostly textual, if they contain links to the original content where the claim was made, it is sometime possible to get to other modalities, such as images, videos or audio, which can be contained in a piece of content (e.g., a social media post). These can be important, because in many cases it is there where the actual claim is being made or the claim requires multiple modalities to be properly interpreted. Multimodal content can also be perceived as more credible by users, has a higher engagement and is increasingly easier to produce (Akhtar et al., 2023). Thus, multimodal approaches continue to be increasingly prevalent as well as important. At the moment, most approaches transcribe the modality to text by either using OCR or image description approaches (Mansour et al., 2022; Pikuliak et al., 2023), however there are already some approaches that process the other modalities directly, be it images (Vo and Lee, 2020) or speech (Ivanov et al., 2024). The future advances will likely lie in advancements of the latter category of approaches.
@@
Adoption of LLMs. Although we observe first approaches using LLMs in check-worthiness detection (Barrón-Cedeño et al., 2024), their potential is still not explored in the surveyed works on the retrieval of fact-checked claims. Their benefit could be in input claim normalization (Sundriyal et al., 2023), using retrieval augmented generation (Huang et al., 2022), in results reranking or in providing summaries of retrieved fact-checks.
@@ 8. Additional credibility signals
8.1. Machine-generated text
The task of machine-generated text detection (also called synthetic or neural text detection or authorship attribution (Uchendu et al., 2023)) has already been researched since about 2018 when the arrival of GPT-1 and transformer architecture have enabled generation of reasonably coherent texts(Uchendu et al., 2023). However, it has been only with the arrival of generative LLMs that the task gained upon practical importance due to increased fidelity and quality of the machine-generated texts in English (Uchendu et al., 2021), but also in many other languages (Macko et al., 2023; Wang et al., 2024b) and due to low costs of such generation. Since then, there has been recognised potential of the machine-generated texts to be misused for influence operations (Goldstein et al., 2023), disinformation (Buchanan et al., 2021; Vykopal et al., 2024; Lucas et al., 2023), spam or unethical authorship (Crothers et al., 2023). Although there have been voices tempering down fears of massive misuse of LLMs for disinformation generation (Simon et al., 2023), actual field studies are missing and the rapid advancement of the new models makes the misuse of LLMs to generate and/or amplify disinformation increasingly easier. Nevertheless, when considering using machine-generated text detection as a credibility signal, it has to be noted that there are also many benign and legitimate uses (e.g., machine translation, use of LLMs to improve stylistics or grammar, etc.), so the context of such a use needs to be taken into account as well.
Due to increased interest in the topic, there have been lately several new datasets released extending the task to a range of languages and domains (Macko et al., 2023; Wang et al., 2024b), often supported by specialized data challenges and tasks, such as SemEval-2024 Task 8 (Wang et al., 2024a). As to the detection methods and models, these range from statistical, e.g., Binoculars (Hans et al., 2024), Fast-DetectGPT (Bao et al., 2024), etc. to fine-tuned language models, such as Longformer (Beltagy et al., 2020; Li et al., 2024a), RoBERTa (Solaiman et al., 2019) or MDeBERTa (Macko et al., 2023), but also increasingly including fine-tuned LLMs or their combinations (Spiegel and Macko, 2024).
@@ 8.2. Text quality
Various NLP techniques have been developed to rate text quality and, by extension, its credibility. Mosquera and Moreda (2021) evaluates text by extracting features like contractions, slang, misspellings, emoticons and readability (using the Readability Index). It also measures information content through entropy and emotional tone using emotional distance. Finally, it applies the Expectation-Maximization (EM) algorithm to cluster texts by informality levels. The tool obtained a F1 score of 60.6%.
In Pitler and Nenkova (2008), the model combines lexical, syntactic, and discourse features to predict human judgments of text readability. It evaluates vocabulary difficulty using unigram language models, syntactic complexity by measuring parse tree height, and discourse relations through annotated markers. Additionally, entity coherence is examined by analyzing semantic continuity between sentences. When all features are used to feed a linear regression, the accuracy on readability results in 88.88% . The model proposed by Mesgar and Strube (2018) uses neural networks to evaluate the quality of the text. It captures semantic connections between adjacent sentences by representing sentences as vectors and identifying the most similar states between them. The model uses Recurrent Neural Networks to account for word context within sentences and a Convolutional Neural Network to identify patterns of coherence across the text. This allows the analysis of sentence-to-sentence transitions to predict the readability with a 97.77% accuracy.
@@ 8.3. References and citations
In Baier et al. (2022), it is discussed how the explicitness of references and the depth of assurance provided by citations influence the perceived credibility of the information with experimental results showing that news articles with explicit citations are perceived as more credible by readers; whereas Hamborg et al. (2017) highlights the importance of citations in presenting diverse perspectives. This is done leveraging the aggregating news from various sources and examining their references and the methodology involves analyzing citation patterns to identify biases and credibility in news articles. References, citations and quotes also can be used in network analysis to detect biases (Aires et al., 2020a). Patricia Aires et al. (2019) used references and citations in link analysis and bias detection, while quotes are discussed in regards to the effectiveness for bias detection.
The datasets used in these studies are obtained from multiple domains, including journalism, social media, and scientific content. In Baier et al. (2022), the dataset comprises articles with varying levels of citation explicitness to experimentally assess reader perceptions of credibility. Hamborg et al. (2017) aggregate news articles from diverse sources, examining their citation patterns to identify biases and provide multiple perspectives. Similarly, works (Aires et al., 2020a; Patricia Aires et al., 2019) derived the datasets from media outlets, focusing on link analysis and citation patterns to detect bias and assess credibility. These datasets reflect a broad range of content types, enhancing the generalizability of the findings across different informational contexts.
@@ 8.4. Clickbaits and title representativeness
To detect clickbait, various NLP methods can be employed, focusing on analyzing lexical and semantic features of headlines. For instance, frequent use of sensationalist language, unresolved pronouns, and forward-referencing structures can indicate clickbait (Zhang and Clough, 2020). Traditional machine learning models, such as Support Vector Machines (SVM) and Naïve Bayes classifiers, are often used to identify these features by assigning probabilities to words and phrases, which are then used to classify headlines as clickbait or non-clickbait (Chen et al., 2015). In this direction, Zhang and Clough (2020) focused on Chinese social media and found significant regional variations in clickbait prevalence. Their studies highlight the effectiveness of SVM and Naïve Bayes classifiers, achieving an F1-measure of 0.834. Combining these approaches with other models like Long Short-Term Memory (LSTM) networks can further enhance classification accuracy, with the SVM model achieving 98.53% accuracy (Winarto et al., 2023).
@@
The datasets used in these studies vary in their scope and origin. For example, one of them focused on Chinese social media platforms like WeChat, where they analyzed regional clickbait patterns using a dataset of social media posts (Zhang and Clough, 2020); whereas the study by Winarto et al. (2023) leveraged YouTube titles in their study, creating a large, labeled dataset of video headlines to train their machine learning models. Additionally, Spezzano et al. (2021) collected news articles with varying degrees of title representativeness, image content, and source bias, analyzing how these elements impacted credibility perceptions. Together, these datasets provide a diverse foundation for training and evaluating models aimed at detecting clickbait across different platforms and content types.

### v2

@@ 7.1. Datasets
Check-worthy claim detection is a popular task within the NLP community thanks to the series of CheckThat! shared tasks organised at CLEF, which led to the release of related datasets. Indeed, check-worthy claim detection is the only subtask that has been proposed at all seven CheckThat! editions (Atanasova et al., 2018, 2019; Barrón-Cedeño et al., 2020; Nakov et al., 2021, 2022a; Alam et al., 2023; Barrón-Cedeño et al., 2024). Through those editions, datasets for training check-worthy claim detection models have been constantly extended to cover additional languages, starting from English and Arabic at CheckThat! 2018 (Atanasova et al., 2018) to Arabic, Dutch, English and Spanish at CheckThat! 2024 (Barrón-Cedeño et al., 2024) and even multimodal and multi-genre data in the 2023 edition (Alam et al., 2023). Beside CheckThat!, additional datasets have been created and made available for research, as shown in Table 8. The sources covered by such datasets are typically social media, news and political debates, i.e. three relevant areas where the presence of disinformation may have detrimental effects on a large audience. Particular relevance was given to COVID-related content during the pandemic (Hadj Ameur and Aliane, 2021; Salek Faramarzi et al., 2023), when the continuous flow of misleading information could negatively affect public health. Although English is the most represented language, datasets for check-worthiness detection have also been developed in several other languages. This can be attributed to the fact that public relevance is both time- and geographically bounded, with claims typically referring to events specific to individual countries or regions.
| Dataset | Language(s) | # Instances | Content type | Classes | ||||||
@@
| Dataset | Language(s) | # Instances | Content type | ||
| # Input claims | # Verified claims | # Pairs | |||
@@ 7.2. Methods and models
To propose methods for check-worthy claim detection, existing works first define how check-worthiness is operationalized. In a recent survey, Panchendrarajan and Zubiaga (2024) identify two main aspects making a claim check-worthy: its verifiability and its priority. The first item refers to the possibility of determining the veracity of a claim, which can be likely supported by evidence. A claim is verifiable when it contains a factual statement that can be checked, which means that personal opinions or events presented as uncertain are excluded. Second, given that verifying all statements about the world is impossible, it is important to prioritize claims which are considered timely, interesting to the general public and whose verification might have a broader impact (Das et al., 2023). In this latter aspect it differs from a related, but a distinct task of claim detection which solely aims to identify what constitutes a claim in a text (either using a binary classification or by identification of spans) (Mittal et al., 2023; Wuehrl et al., 2023; Gupta et al., 2021b).
@@
Given this need for prioritization, check-worthiness detection has been cast as a ranking problem since the first editions of the CheckThat! Lab within the CLEF Evaluation initiative232323https://www.clef-initiative.eu/, which made the task well-known within the NLP community, although some related works had already been presented before (Hassan et al., 2015). In particular, given a political debate, the first CheckThat! task was aimed at predicting which claims should be prioritized for fact checking. This is reflected in the evaluation approach proposed for the CheckThat! series, which has become the de facto standard in check-worthiness detection: systems should output the list of input claims ranked by check-worthiness, which is usually evaluated using Mean Average Precision (MAP), reciprocal rank and for . More recently, check-worthiness has been, however, evaluated using F1 as a binary classification task rather than ranking (Alam et al., 2023; Barrón-Cedeño et al., 2024).
@@
Regarding methods, in the CheckThat! editions up to 2023 state-of-the-art results for check-worthy claim detection were reached by methods that rely on fine-tuned transformer-based methods such as BERT, RoBERTa, DistilBERT, (Williams et al., 2021; Frick et al., 2023; Sawinski et al., 2023) and language-specific variants, (Alam et al., 2023), often combined with data manipulation or model ensembling strategies. For example, the best results on the English portion of the CheckThat 2022 dataset were achieved by a RoBERTa model that leveraged a back-translation-driven data augmentation process (Savchev, 2022). The 2024 edition, instead, has seen an increased interest in using also generative LLMs for the task such as LLama 2 and 3, GPT3.4 and 5, Mixtral and Mistral (Barrón-Cedeño et al., 2024). For instance, the best performing system on the English language dataset relied on fine-tuned Llama2 7b on the original training data, using prompts generated by ChatGPT (Li et al., 2024b). Nevertheless, despite the notable progress in data and methods for check-worthy claim detection, there is currently a lack of an extended coverage across languages and topics.
@@ 7.3. Tools and services
ClaimBuster (Hassan et al., 2017) was the first end-to-end system for computer-assisted fact-checking trained on a human-labeled dataset of check-worthy factual claims from the U.S. general election debate transcripts. The first component of the pipeline is a detector of check-worthy factual claims which, given a sentence, first labels it as being ‘non-factual’, ‘factual and unimportant’ or ‘factual and check-worthy’. In case of the latter, a ranking score is assigned based on SVM decision function. Patwari et al. (2017) present Tathya, a tool focusing only on check-worthiness detection, which compared to ClaimBuster can yield a significant performance improvement, particularly on recall. It is based on a multi-classifier system using features such as topics, entity history and PoS tuples.
ClaimRank (Jaradat et al., 2018) performs check-worthy claim detection and supports English and Arabic texts. Its strength is that it was trained on actual annotations from nine reputable fact-checking organizations, therefore mimicking their real strategy for claim selection. The ranking is based on a number of lexical, structural and semantic features, used to train a neural network with two hidden layers as proposed by Gencheva et al. (2017). Another system, focusing specifically on tweets in Arabic, is Tahaqqaq (Sheikh Ali et al., 2023), which includes the possibility to identify check-worthy claims, estimate the user credibility in terms of spreading fake news, and find authoritative accounts.
@@
Concerning retrieval of previously fact-checked claims, it is supported by some of the end-to-end verification systems mentioned above, namely ClaimBuster (Hassan et al., 2017), Tahaqqaq (Sheikh Ali et al., 2023) or BRENDA (Botnevik et al., 2020). Besides these, Google Fact-Check Explorer242424https://toolbox.google.com/factcheck/explorer is often used to perform the task since it indexes a large corpus of fact-checks. Other specialized tools include Fact-Check Finder252525https://fact-check-finder.kinit.sk built on models developed in (Pikuliak et al., 2023).
@@ 7.4. Discussion
Multimodality. Although most available datasets are mostly textual, if they contain links to the original content where the claim was made, it is sometime possible to get to other modalities, such as images, videos or audio, which can be contained in a piece of content (e.g., a social media post). These can be important, because in many cases it is there where the actual claim is being made or the claim requires multiple modalities to be properly interpreted. Multimodal content can also be perceived as more credible by users, has a higher engagement and is increasingly easier to produce (Akhtar et al., 2023). Thus, multimodal approaches continue to grow in both prevalence and importance. At the moment, most approaches transcribe the modality to text by either using OCR or image description approaches (Mansour et al., 2022; Pikuliak et al., 2023), however, there are already some approaches that process the other modalities directly, be it images (Vo and Lee, 2020) or speech (Ivanov et al., 2024). The future advances will likely lie in advancements of the latter category of approaches.
@@
Adoption of LLMs. Although we observe some approaches using LLMs in check-worthiness detection (Barrón-Cedeño et al., 2024), their potential for the retrieval of fact-checked claims begins only now to be more systematically explored. Most recently, LLMs have been used and evaluated as text embedding models as well as rerankers of retrieved fact-checked claims in multilingual and crosslingual settings in (Ramponi et al., 2025), outperforming the fine-tuned smaller models. They have been also employed in zero- and few-shot settings as binary classifiers filtering out irrelevant retrieved results using a range of prompting strategies (Vykopal et al., 2025b; Pisarevskaya and Zubiaga, 2025); these recent works showed that while useful also in this setup, no single prompting strategy proved as the best overall and the performance of the current LLMs is lower for low-resource languages compared to high-resource ones. The use of LLMs (as text embedders, filters or rerankers) have been further explored in system papers submitted to the recent SemEval-2025 Task 7 as summarised in (Peng et al., 2025). Finally, the benefit of LLMs for both check-worthiness detection or retrieval of previously fact-checked claims lies in input claim normalization (Sundriyal et al., 2023), using retrieval augmented generation (Huang et al., 2022), or in providing summaries of the check-worthy content or of the retrieved fact-checks (Vykopal et al., 2025a).
@@ 8.1. Text quality
Various NLP techniques have been developed to rate text quality and, by extension, its credibility. Mosquera and Moreda (2021) evaluated text by extracting features like contractions, slang, misspellings, emoticons and readability (using the Readability Index). They also measured information content through entropy and emotional tone using emotional distance and, finally, applied the Expectation-Maximization (EM) algorithm to cluster texts by informality levels. The tool obtained a F1 score of 60.6%. In Pitler and Nenkova (2008), the proposed model combines lexical, syntactic, and discourse features to predict human judgments of text readability. It evaluates vocabulary difficulty using unigram language models, syntactic complexity by measuring parse tree height, and discourse relations through annotated markers. Additionally, entity coherence is examined by analysing semantic continuity between sentences. When all features are used to feed a linear regression, the accuracy on readability results in 88.88%. The model proposed by Mesgar and Strube (2018) uses neural networks to evaluate the quality of the text. It captures semantic connections between adjacent sentences by representing sentences as vectors and identifying the most similar states between them. The model uses Recurrent Neural Networks to account for word context within sentences and a Convolutional Neural Network to identify patterns of coherence across the text. This allows the analysis of sentence-to-sentence transitions to predict the readability with a 97.77% accuracy.
@@ 8.2. References and citations
In Baier et al. (2022), it is discussed how the explicitness of references and the depth of assurance provided by citations influence the perceived credibility of the information with experimental results showing that news articles with explicit citations are perceived as more credible by readers. Similarly, Hamborg et al. (2017) highlights the importance of citations in presenting diverse perspectives. Besides overall credibility, references, citations and quotes also can be used in network analysis to detect biases (Aires et al., 2020a). Patricia Aires et al. (2019) used references and citations in link analysis and bias detection, while quotes are discussed in regards to the effectiveness for bias detection.
The datasets used in these studies are obtained from multiple domains, including journalism, social media, and scientific content. In Baier et al. (2022), the dataset comprises articles with varying levels of citation explicitness to experimentally assess reader perceptions of credibility. Hamborg et al. (2017) aggregated news articles from diverse sources and examined their references and citation patterns. Similarly, works (Aires et al., 2020a; Patricia Aires et al., 2019) derived the datasets from media outlets, focusing on link analysis and citation patterns to detect bias and assess credibility. These datasets reflect a broad range of content types, enhancing the generalizability of the findings across different informational contexts.
@@ 8.3. Clickbaits and title representativeness
To detect clickbait, various NLP methods can be employed, focusing on analysing lexical and semantic features of headlines. For instance, frequent use of sensationalist language, unresolved pronouns, and forward-referencing structures can indicate clickbait (Zhang and Clough, 2020). Traditional machine learning models, such as Support Vector Machines (SVM) and Naïve Bayes classifiers, are often used to identify these features by assigning probabilities to words and phrases, which are then used to classify headlines as clickbait or non-clickbait (Chen et al., 2015). In this direction, Zhang and Clough (2020) focused on Chinese social media and found significant regional variations in clickbait prevalence. Their studies highlight the effectiveness of SVM and Naïve Bayes classifiers, achieving an F1-measure of 0.834. Combining these approaches with other models like Long Short-Term Memory (LSTM) networks can further enhance classification accuracy, with the SVM model achieving 98.53% accuracy (Winarto et al., 2023).
@@
The datasets used in these studies vary in their scope and origin. For example, one of them focused on Chinese social media platforms like WeChat, where they analysed regional clickbait patterns using a dataset of social media posts (Zhang and Clough, 2020); whereas the study by Winarto et al. (2023) leveraged YouTube titles in their study, creating a large, labeled dataset of video headlines to train their machine learning models. Additionally, Spezzano et al. (2021) collected news articles with varying degrees of title representativeness, image content, and source bias, analysing how these elements impacted credibility perceptions. Together, these datasets provide a diverse foundation for training and evaluating models aimed at detecting clickbait across different platforms and content types.